Qx=200-px+2py+.03y where px=10 py=15 y=1000.

Find the PED.


1
Expert's answer
2021-11-09T10:47:59-0500

"PED=\delta Q/\delta P*P/Q"

"PED=-1*P\/Q"

"Q=200-10+2(15)+0.03(1000)"

"Q=200-10+30+30"

"Q=250"

"PED=-1*10\/250"

"PED=-0.04"